Write equation of the line that is the perpendicular bisector of the given segment. A(-4,-7),B(0,1). Write equation in Slope intercept Form.
Answer by josgarithmetic(39616)

How to find the line asked for :

Find slope of AB.
Find negative reciprocal of that slope and call the result m.
Find midpoint of AB and this will be a specific point, (u,v).
Use point-slope form to formulate your line using slope m and using the point (u,v).
y-v=m(x-u);
Perform steps to transform the equation into slope-intercept form, y=mx+b.

You try the steps.